OpenStudy (anonymous):

x^2 – 8x + 15 factor completely

OpenStudy (anonymous):

\[x^2 – 8x + 15\] \[8=3+5\]\[15=3*5\] \[x^2-8x+15=(x-3)(x-5)\]

OpenStudy (chmvijay):

(x-3)(x-5)

OpenStudy (anonymous):

x^2-5x-3x+15= x(x-5)-3(x-5)= (x-5)(x-3)